Draw all possible simples of size 2 with replacement from 3,5,7,8,9.And verify the rezult.


1
Expert's answer
2021-02-24T05:58:44-0500

There are "n^2=5^2=25" samples of size two which can be drawn with replacement: 


"\begin{matrix}
 (3,3) & (3,5) & (3, 7) & (3, 8) & (3, 9) \\
 (5,3) & (5,5) & (5, 7) & (5, 8) & (5, 9) \\
 (7,3) & (7,5) & (7, 7) & (7, 8) & (7, 9) \\
 (8,3) & (8,5) & (8, 7) & (8, 8) & (8, 9) \\
 (9,3) & (9,5) & (9, 7) & (9, 8) & (9, 9) \\
\end{matrix}"

There are 25 samples of size 2 with replacement from 3,5,7,8,9.